OpenStudy (anonymous):

Factor: 27x4-12(xcubed)-36/(xcubed) please and thank you :)

OpenStudy (anonymous):

\[27x ^{4}-12\sqrt[3]{x}-36/\sqrt[3]{x}\]

OpenStudy (anonymous):

Is that \(x^3\) or \(\sqrt[3]{x}\) ??

OpenStudy (anonymous):

\[\sqrt[3]{x} :)\]

OpenStudy (anonymous):

let \[\sqrt[3]{x}\] be a

OpenStudy (anonymous):

and then how do I factor the 36/a along with the rest of the equation?

OpenStudy (anonymous):

\[a^{12}-12a-\frac{ 36 }{ a }\]

OpenStudy (anonymous):

\[a^{13}-12a^{2}-36\]

OpenStudy (anonymous):

what happened to the 27x^4?

OpenStudy (anonymous):

\[x^{4}= \sqrt[3]{x}^{12}\]

OpenStudy (anonymous):

after y^13-12y^2-36 what should i do?

OpenStudy (anonymous):

try to factor out then for the roots you find put \[\sqrt[3]{x} \] instead of y

OpenStudy (anonymous):

and then what? D:

OpenStudy (anonymous):

sorry i don't really know anything

OpenStudy (anonymous):

did you find the roots

OpenStudy (anonymous):

im unsure of what that means

OpenStudy (anonymous):

for example a^2-2a-1=(a-1)(a-1) then the root is a=1